1. J.R.R. Tolkien wrote in his epie novel The Lord of the Rings that the One Ring ..did not seem always of the same size or weight; it shrank and expanded in an odd way... For instance, in Peter Jacksons film adaptation of The Fellouship of the Ring, Isildur cut the One Ring from the hand of the Dark Lord Sauron during the final battle on the slopes of Mount Doom, and once the ring was removed from the severed finger of Sauron, it contracted to the size of a large male finger, which is about 25 mm in diameter. Lets assume the ring changes size according to thermal expansion and was made of a material with a coefficient of thermal expansion of 40 x 10-5 K-. Suppose, too, that Suaron, who is often pictured to be standing in fire or lava, has a temperature of 1100° C. If the ring cools to a temperature of 30° C when it contracts, what is the diameter of Saurons finger if we assume the ring fit snuggly before it was removed?
Answer #1

Given data

alpha=40*10^-5

delta T=-1100+30

=-1070

initial diameter=25mm

final diameter=25( 1-40*10^-5*1070)

=14.3mm